# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 2. A boundary value problem for the Dirac operator in odd dimensions
Since $n$ is odd, the restriction $\mathcal{S}|_{\Sigma}$ can be canonically identified with the spinor bundle of $\Sigma$. The field of involutions $i\gamma(\nu)$ can be used to decompose
$\mathcal{S}|_{\Sigma}=\mathcal{S}_{\Sigma}^{+}\oplus\mathcal{S}_{\Sigma}^{-}$
where $\mathcal{S}_{\Sigma}^{\pm}$ are the eigensubbundles corresponding to the eigenvalues ±$\pm 1$. Let $\mathcal{S}_{0}$ denote the trivial bundle over $S^{n-1}$ with fiber $S_{0}$, and let $\mathcal{S}_{0}^{*}$ denote the dual bundle over $S^{n-1}$ with fiber $S_{0}^{*}$. With this understood, we may write $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$. We may decompose the bundle $\mathcal{S}_{0}^{*}$ as
@@
where, for each point $v\in S^{n-1}$, $\mathcal{S}_{0}^{*,\pm}$ denote the eigensubbundles of $i\gamma(v)$ corresponding to the eigenvalues ±$\pm 1$. This gives a decomposition
(2.5) $N^{*}\mathcal{S}^{*}=N^{*}\mathcal{S}_{0}^{*,+}\oplus N^{*}\mathcal{S}_{0}^{*,-},$
where $N^{*}\mathcal{S}_{0}^{*,\pm}$ are the eigensubbundles of $i\gamma_{0}^{*}(N)$ corresponding to the eigenvalues ±$\pm 1$. We then have
@@
We next consider the twisted Dirac operators
@@
$\mathcal{D}^{\Sigma}_{-}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-}).$
Then the operator
(2.6) $\mathcal{D}^{\Sigma}=\begin{pmatrix}0&\mathcal{D}^{\Sigma}_{-}\\ \mathcal{D}^{\Sigma}_{+}&0\end{pmatrix}$
maps sections of $\mathcal{F}^{+}$ to sections of $\mathcal{F}^{-}$.
In the following, $\hat{\mathsf{A}}(\Sigma)$ denotes the $\hat{\mathsf{A}}$-form of the tangent bundle of $\Sigma$ and $\mathsf{ch}(\mathcal{S}_{0}^{*,\pm})$ the Chern character form of the bundle $\mathcal{S}_{0}^{*,\pm}$. The lower index $k$ in $\hat{\mathsf{A}}(\Sigma)_{k}$ indicates the homogeneous part of degree $k$ and similarly for the Chern character. By the Atiyah-Singer index theorem, we have
@@
$\mathsf{ind}(\mathcal{D}^{\Sigma})$ $=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$
@@
Now, we observe that $\mathcal{D}$ and $\mathcal{D}^{\Sigma}$ are formally selfadjoint, $\mathcal{D}^{\Sigma}$ is an adapted boundary operator for $\mathcal{D}$, the involution $i\gamma(\nu)\otimes 1$ preserves the splitting (2.3) and anticommutes with $\mathcal{D}^{\Sigma}$. Therefore, the holographic index theorem from Appendix B in [BBHW23] applies and tells us that the operator
is Fredholm and
$\mathsf{ind}(\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E}))=\tfrac{1}{2}\,\mathsf{ind}(\mathcal{D}^{\Sigma})=\mathsf{deg}(N).\qed$
@@ 3. Spin geometry for initial data sets
Let $M$ be a compact spin manifold of dimension $n\geq 2$ with boundary $\partial M=\Sigma$. Let $g$ be a Riemannian metric on $M$, let $q$ be a symmetric $(0,2)$-tensor field on $M$ and let $N\colon\Sigma\to S^{n-1}$ be a smooth map. Let $\chi$ be defined as in (2.2). We will denote the selfadjoint endomorphism field corresponding to $q$ also by $q\colon TM\to TM$, i.e.
$q(X,Y)=g(q(X),Y).$
Furthermore, we put
@@
(3.2) $J:={\mathrm{div}}(q)-\nabla\mathrm{tr}(q)$
where $R$ is the scalar curvature of $g$. We use $q$ to modify the connections on $\mathcal{S}$ and on $\mathcal{E}$.
@@
(3.3) $\nabla^{\mathcal{S},q}:=\nabla^{\mathcal{S}}+\frac{i}{2}\gamma\circ q$
on $\mathcal{S}$. We next introduce the homomorphism field $Q\colon\mathcal{E}\to T^{*}M\otimes\mathcal{E}$ which maps spinors to spinor-valued $1$-forms by setting
(3.4) $(Qs)(X):=(\gamma(q(X))\otimes 1)s$
@@
(3.5) $\nabla^{\mathcal{E},q}:=\nabla^{\mathcal{E}}+\tfrac{i}{2}Q.$
Note that the connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ with the flat connection on the trivial bundle over $M$ with fiber $S_{0}$.
@@
Denote the (pointwise) adjoint of $Q$ by $Q^{*}\colon T^{*}M\otimes\mathcal{E}\to\mathcal{E}$. Then we have
@@
    $Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q=\gamma(\mathrm{div}(q))\otimes 1$.
@@
Let $p\in M$ and let $e_{1},\dots,e_{n}$ be an orthonormal basis of $T_{p}M$. Denote the dual basis of $T^{*}_{p}M$ by $e^{1},\dots,e^{n}$. Then we have for $s\in\mathcal{E}|_{p}$ and $\psi\in T^{*}_{p}M\otimes\mathcal{E}|_{p}$:
$\langle Q^{*}\psi,s\rangle$ $=\langle\psi,Qs\rangle=\Big{\langle}\psi,\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{\rangle}$
$=\sum_{j}\langle\psi(e_{j}),(\gamma(q(e_{j}))\otimes 1)s\rangle$
$=-\sum_{j}\langle(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}),s\rangle$
and hence
$Q^{*}\psi=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}).$
@@
$Q^{*}Qs$ $=Q^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
$=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)(\gamma(q(e_{j}))\otimes 1)s$
@@
This proves (a). To show (b), extend the orthonormal basis to a neighborhood of $p$ such that it is synchronous at $p$. Let $s$ be a smooth section of $\mathcal{E}$, defined near $p$. Then we compute at $p$:
$(\nabla^{\mathcal{E}})^{*}Qs$ $=(\nabla^{\mathcal{E}})^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
$=-\sum_{j}(\gamma((D_{e_{j}}q)(e_{j}))\otimes 1)s-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\nabla_{e_{j}}^{\mathcal{E}}s$
$=-(\gamma(\mathrm{div}(q))\otimes 1)s+Q^{*}\nabla^{\mathcal{E}}s.\qed$
The well-known Weitzenböck formula gives
$\mathcal{D}^{2}=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}+\tfrac{1}{4}R.$
We find a modification, taking $q$ into account.
@@
$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))=(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}+\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.$
@@
$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}\gamma(\mathrm{div}(q))\otimes 1+\tfrac{1}{4}|q|^{2}.$
@@
$=\mathcal{D}^{2}-\tfrac{i}{2}\gamma(\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}\mathrm{tr}(q)^{2}.$
@@
$=\tfrac{1}{4}R+\tfrac{i}{2}\gamma(\mathrm{div}(q)-\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}(\mathrm{tr}(q)^{2}-|q|^{2})$
$=\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.\qed$
Again we denote $\partial M=\Sigma$ and the unit outward normal along $\Sigma$ by $\nu$. Let $H$ denote the (unnormalized) mean curvature. The sign convention is such that the boundary of a Euclidean ball has positive mean curvature.
We define an operator $\mathcal{D}^{\Sigma}_{0}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
$\mathcal{D}^{\Sigma}_{0}=-\sum_{j=1}^{n-1}(\gamma(\nu)\otimes 1)(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}-\tfrac{1}{2}H,$
where $e_{1},\ldots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$ and $\nabla^{\mathcal{E}}$ denotes the connection on $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ defined above. By Proposition 2.2 in [B96], $\mathcal{D}^{\Sigma}_{0}$ can be identified with a twisted Dirac operator acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$, where the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ carries the flat connection. Note that $\mathcal{D}^{\Sigma}_{0}$ differs from the operator $\mathcal{D}^{\Sigma}$ defined in (2.6). In the definition of $\mathcal{D}^{\Sigma}$, we used a nonflat connection on the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ that preserves the splittting (2.5). The operators $\mathcal{D}^{\Sigma}$ and $\mathcal{D}^{\Sigma}_{0}$ both anticommute with $\gamma(\nu)\otimes 1$.
@@
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}$ $+\int_{M}|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$=\int_{\Sigma}\langle(-\mathcal{D}^{\Sigma}_{0}-\tfrac{1}{2}H+\tfrac{i}{2}(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1))s,s\rangle.$
@@
$(\gamma(\nu)\otimes 1)\mathcal{D}s=-\mathcal{D}^{\Sigma}_{0}s-\nabla_{\nu}^{\mathcal{E}}s-\tfrac{1}{2}Hs$
along $\Sigma$. Using this and integrating by parts we find
@@
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\langle(\gamma(\nu)\otimes 1)(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\big{\langle}\big{(}-\mathcal{D}^{\Sigma}_{0}-\nabla_{\nu}^{\mathcal{E}}-\tfrac{1}{2}H-\tfrac{i}{2}\mathrm{tr}(q)(\gamma(\nu)\otimes 1)\big{)}s,s\big{\rangle}.$
@@
$=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle(\nabla_{\nu}^{\mathcal{E}}+\tfrac{i}{2}\gamma(q(\nu))\otimes 1)s,s\rangle.$
Inserting these two equations into the modified Weitzenböck formula concludes the proof. ∎
Lemma 3.5.
Along $\Sigma$ we have
$\mathcal{D}^{\Sigma}_{0}\chi+\chi\mathcal{D}^{\Sigma}_{0}=\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j}))$
Proof.
Recall that $\mathcal{D}^{\Sigma}_{0}$ anticommutes with $\gamma(\nu)\otimes 1$. Hence, we compute for a smooth section $s$ of $\mathcal{E}|_{\Sigma}$:
$\mathcal{D}^{\Sigma}_{0}\chi s$ $=-\mathcal{D}^{\Sigma}_{0}(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$=(\gamma(\nu)\otimes 1)\mathcal{D}^{\Sigma}_{0}(1\otimes\gamma_{0}^{*}(N))s$
$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}((1\otimes\gamma_{0}^{*}(N))s)-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\nabla_{e_{j}}^{\mathcal{E}}s-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$\quad+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=-\chi\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s.\qed$
@@
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{M}$ $|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$\leq\tfrac{1}{2}\int_{\Sigma}(\|dN\|_{\mathrm{tr}}-H+|q(\nu,\nu)-\mathrm{tr}(q)|)|s|^{2}.$
@@
$2\,\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s,s\rangle+\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}s,\chi s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s+\chi\mathcal{D}^{\Sigma}_{0}s,s\rangle$
@@
$2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|=\Big{|}\Big{\langle}\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big{\rangle}\Big{|}$
@@
(3.6) $2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
@@
$\langle(\gamma(X)\otimes 1)s,s\rangle$ $=\varepsilon\langle(\gamma(X)\otimes 1)\chi s,s\rangle$
$=-\varepsilon\langle\chi(\gamma(X)\otimes 1)s,s\rangle$
$=-\varepsilon\langle(\gamma(X)\otimes 1)s,\chi s\rangle$
$=-\langle(\gamma(X)\otimes 1)s,s\rangle,$
hence
$\langle(\gamma(X)\otimes 1)s,s\rangle=0.$
@@
$\langle(\gamma(q(\nu))\otimes 1)s,s\rangle=q(\nu,\nu)\,\langle(\gamma(\nu)\otimes 1)s,s\rangle.$

### v4

@@ 2. A boundary value problem for the Dirac operator in odd dimensions
Note that $\mathcal{D}$ and $\mathcal{A}$ are formally self-adjoint, and $\mathcal{A}$ is an adapted boundary operator for $\mathcal{D}$. It follows from Lemma 2.3 that $\mathcal{A}$ is an odd operator in the sense that it interchanges the bundles $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$. Moreover, the involution $i\gamma(\nu)\otimes\textup{id}$ preserves the splitting (2.2) and anticommutes with $\mathcal{A}$. Therefore, the holographic index theorem [Baer-Brendle-Hanke-Wang]*Theorem B.1 applies and tells us that the operator
is Fredholm and
$\mathsf{ind}(\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E}))=\tfrac{1}{2}\,\mathsf{ind}(\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-})).$
It remains to compute the index of the operator $\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-})$. Since $n$ is odd, the restriction $\mathcal{S}|_{\Sigma}$ can be canonically identified with the spinor bundle of $\Sigma$. The field of involutions $i\gamma(\nu)$ gives a decomposition $\mathcal{S}|_{\Sigma}=\mathcal{S}_{\Sigma}^{+}\oplus\mathcal{S}_{\Sigma}^{-}$, where $\mathcal{S}_{\Sigma}^{\pm}$ are the eigensubbundles corresponding to the eigenvalues $\pm 1$. Moreover, we may decompose the bundle $\mathcal{S}_{0}^{*}$ as
@@
where, for each point $v\in S^{n-1}$, $\mathcal{S}_{0}^{*,\pm}$ denote the eigensubbundles of $i\gamma(v)$ corresponding to the eigenvalues $\pm 1$. This gives a decomposition
$N^{*}\mathcal{S}_{0}^{*}=N^{*}\mathcal{S}_{0}^{*,+}\oplus N^{*}\mathcal{S}_{0}^{*,-},$
where $N^{*}\mathcal{S}_{0}^{*,\pm}$ are the eigensubbundles of $i\gamma_{0}^{*}(N)$ corresponding to the eigenvalues $\pm 1$. We then have
@@
Recall that $\mathcal{A}$ maps sections of
$\mathcal{F}^{+}=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})$
to sections of $\mathcal{F}^{-}=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+})$. Since the involution $i\gamma(\nu)\otimes\textup{id}$ preserves the splitting (2.2) and anticommutes with $\mathcal{A}$, it follows that $\mathcal{A}$ maps sections of $\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+}$ to sections of $\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+}$, and $\mathcal{A}$ maps sections of $\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-}$ to sections of $\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-}$. Therefore, the operator $\mathcal{A}\colon\mathcal{F}^{+}\to\mathcal{F}^{-}$ can be written in the form
$\begin{bmatrix}0&\mathcal{D}^{\Sigma}_{-}\\ \mathcal{D}^{\Sigma}_{+}&0\end{bmatrix}$,
where
@@
$\mathcal{D}^{\Sigma}_{-}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-})$
are Dirac-type operators. Thus,
$\mathsf{ind}(\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-}))=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-}).$
We may compute $\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})$ and $\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$ using the Atiyah-Singer index theorem. To that end, we denote by $\hat{\mathsf{A}}(\Sigma)$ the $\hat{\mathsf{A}}-$form of the tangent bundle of $\Sigma$, and by $\mathsf{ch}(\mathcal{S}_{0}^{*,\pm})$ the Chern character form of the bundle $\mathcal{S}_{0}^{*,\pm}$. The lower index $k$ in $\hat{\mathsf{A}}(\Sigma)_{k}$ indicates the homogeneous part of degree $k$. We use the analogous notation for the homogeneous parts of the Chern character. The Atiyah-Singer index theorem gives
@@
$\mathsf{ind}(\mathcal{A}\colon$ $H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-}))$
$=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$
@@
Putting these facts together, the assertion follows. ∎
@@ 3. Spin geometry for initial data sets
Let $M$ be a compact spin manifold of dimension $n\geq 2$ with boundary $\partial M=\Sigma$. Let $g$ be a Riemannian metric on $M$, let $q$ be a symmetric $(0,2)$-tensor field on $M$ and let $N\colon\Sigma\to S^{n-1}$ be a smooth map. Let $\chi$ be defined as in (2.1). We will denote the self-adjoint endomorphism field corresponding to $q$ also by $q\colon TM\to TM$, i.e. $q(X,Y)=g(q(X),Y)$. Furthermore, we put
@@
(3.2) $J:=\text{\rm div}(q)-\nabla\mathrm{tr}(q)$
where $R$ is the scalar curvature of $g$. In the first step, we use $q$ to define modified connections on $\mathcal{S}$ and on $\mathcal{E}$.
@@
$\nabla_{X}^{\mathcal{S},q}:=\nabla_{X}^{\mathcal{S}}+\tfrac{i}{2}\gamma(q(X))$
on $\mathcal{S}$. We next define a homomorphism field $Q\colon\mathcal{E}\to T^{*}M\otimes\mathcal{E}$ by setting
$(Qs)(X):=(\gamma(q(X))\otimes\textup{id})s$
@@
$\nabla_{X}^{\mathcal{E},q}s:=\nabla_{X}^{\mathcal{E}}s+\tfrac{i}{2}(Qs)(X).$
Note that the connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ with the flat connection on the trivial bundle over $M$ with fiber $S_{0}^{*}$.
@@
Let $Q^{*}\colon T^{*}M\otimes\mathcal{E}\to\mathcal{E}$ denote the pointwise adjoint of $Q$. Then the following statements hold:
@@
    $Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q=\gamma(\text{\rm div}(q))\otimes\textup{id}$.
@@
Let $p\in M$ and let $e_{1},\dots,e_{n}$ be an orthonormal basis of $T_{p}M$. We denote the dual basis of $T^{*}_{p}M$ by $e^{1},\dots,e^{n}$. For $s\in\mathcal{E}|_{p}$ and $\psi\in T^{*}_{p}M\otimes\mathcal{E}|_{p}$, we compute
$\langle Q^{*}\psi,s\rangle$ $=\langle\psi,Qs\rangle=\Big\langle\psi,\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes\textup{id})s\Big\rangle$
$=\sum_{j}\langle\psi(e_{j}),(\gamma(q(e_{j}))\otimes\textup{id})s\rangle$
$=-\sum_{j}\langle(\gamma(q(e_{j}))\otimes\textup{id})\psi(e_{j}),s\rangle.$
@@
$Q^{*}\psi=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})\psi(e_{j}).$
Consequently,
$Q^{*}Qs$ $=Q^{*}\Big(\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes\textup{id})s\Big)$
$=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})(\gamma(q(e_{j}))\otimes\textup{id})s$
@@
This proves (a). To show (b), we extend the orthonormal basis to a neighborhood of $p$ such that $\nabla e_{j}=0$ at $p$ for $j=1,\ldots,n$. Let $s$ be a smooth section of $\mathcal{E}$, defined near pp. Then we compute at $p$:
$(\nabla^{\mathcal{E}})^{*}Qs$ $=-\sum_{j}\nabla_{e_{j}}^{\mathcal{E}}\big((\gamma(q(e_{j}))\otimes\textup{id})s\big)$
$=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})\nabla_{e_{j}}^{\mathcal{E}}s-\sum_{j}(\gamma((D_{e_{j}}q)(e_{j}))\otimes\textup{id})s$
$=Q^{*}\nabla^{\mathcal{E}}s-(\gamma(\text{\rm div}(q))\otimes\textup{id})s.\qed$
The well-known Weitzenböck formula gives $\mathcal{D}^{2}=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}+\tfrac{1}{4}R$. We find a modification, taking $q$ into account.
@@
$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))=(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}+\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes\textup{id}.$
@@
$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}\gamma(\text{\rm div}(q))\otimes\textup{id}+\tfrac{1}{4}|q|^{2}.$
@@
$=\mathcal{D}^{2}-\tfrac{i}{2}\gamma(\nabla\mathrm{tr}(q))\otimes\textup{id}+\tfrac{1}{4}\mathrm{tr}(q)^{2}.$
@@
$=\tfrac{1}{4}R+\tfrac{i}{2}\gamma(\text{\rm div}(q)-\nabla\mathrm{tr}(q))\otimes\textup{id}+\tfrac{1}{4}(\mathrm{tr}(q)^{2}-|q|^{2})$
$=\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes\textup{id}.\qed$
@@
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}$ $+\int_{M}|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
$=\int_{\Sigma}\langle(\mathcal{D}^{\Sigma}-\tfrac{1}{2}H+\tfrac{i}{2}(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes\textup{id}))s,s\rangle.$
@@
$(\gamma(\nu)\otimes\textup{id})\mathcal{D}s=\mathcal{D}^{\Sigma}s-\nabla_{\nu}^{\mathcal{E}}s-\tfrac{1}{2}Hs$
along $\Sigma$. Using this identity and integration by parts we find
@@
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\langle(\gamma(\nu)\otimes\textup{id})(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\big\langle\big(\mathcal{D}^{\Sigma}-\nabla_{\nu}^{\mathcal{E}}-\tfrac{1}{2}H-\tfrac{i}{2}\mathrm{tr}(q)(\gamma(\nu)\otimes\textup{id})\big)s,s\big\rangle.$
@@
$=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle(\nabla_{\nu}^{\mathcal{E}}+\tfrac{i}{2}\gamma(q(\nu))\otimes\textup{id})s,s\rangle.$
Substituting these equations into the modified Weitzenböck formula concludes the proof. ∎
@@
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{M}$ $|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
$\leq\tfrac{1}{2}\int_{\Sigma}(|q(\nu,\nu)-\mathrm{tr}(q)|+\|dN\|_{\mathrm{tr}}-H)|s|^{2}.$
@@
$2\,\langle\mathcal{D}^{\Sigma}s,s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}\chi s,s\rangle+\varepsilon\,\langle\mathcal{D}^{\Sigma}s,\chi s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}\chi s+\chi\mathcal{D}^{\Sigma}s,s\rangle$
@@
$2\,|\langle\mathcal{D}^{\Sigma}s,s\rangle|=\Big|\Big\langle\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big\rangle\Big|$
@@
(3.3) $2\,|\langle\mathcal{D}^{\Sigma}s,s\rangle|\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
@@
$2\,\langle(\gamma(X)\otimes\textup{id})s,s\rangle$ $=\varepsilon\,\langle(\gamma(X)\otimes\textup{id})\chi s,s\rangle+\varepsilon\,\langle(\gamma(X)\otimes\textup{id})s,\chi s\rangle$
$=\varepsilon\,\langle(\gamma(X)\otimes\textup{id})\chi s+\chi(\gamma(X)\otimes\textup{id})s,s\rangle$
$=0.$
@@
$\langle(\gamma(q(\nu))\otimes\textup{id})s,s\rangle=q(\nu,\nu)\,\langle(\gamma(\nu)\otimes\textup{id})s,s\rangle.$